OpenStudy (unklerhaukus):

Find \(F(x)\) \[F(x)=\int f(x)\text dx\] \[f(x)=0\]

OpenStudy (unklerhaukus):

this is not ment to be difficult

OpenStudy (zepp):

F(x) = 0! :D \(F(x)=\int(0*dx)\)=0! :D

OpenStudy (unklerhaukus):

zero is not the correct answer sorry.

OpenStudy (zepp):

Was just kidding :(

OpenStudy (unklerhaukus):

nice try

OpenStudy (unklerhaukus):

you have forgotten one thing only

OpenStudy (anonymous):

the answer is just a constant

OpenStudy (unklerhaukus):

so the answer could be any real number?

OpenStudy (anonymous):

yes sir

OpenStudy (unklerhaukus):

\[F(x)=c\qquad;\qquad c\in\mathbb R\]